8. (12 pts.) Find all the relative extreme points of the function f(x,y) = y* – 32y + - x?.
Answer #1


Ans: 4 4 x 3 - 12 Jen,y) = yt. 329 find critical point using In = ty zo fr 3 ay tuis mi) by = yt_s2y +9??) In ty = 443-32 In